# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ Conclusion.
Let $Y_{LB}>0$, $Y_{UB}>0$ be bounds for $y\in[Y_{LB},Y_{UB}]$, and $x^{\max}>0$ for $z\in[0,x^{\max}]$. Let $c(x)$ be a continuously differentiable and strongly convex cost function with $\left|c^{\prime}(x)\right|\leq M_{c^{\prime}}$ for all $x\in[0,x^{\max}]$. Assume $\alpha,\mu>0$, and $\gamma\in[0,1]$.
Define $K(\sigma)\coloneqq\frac{\mu(1+\alpha\sigma)}{1+\alpha}$, and let
@@
$\left|\phi^{1}(z_{1},y_{1},\sigma_{1})-\phi^{1}(z_{2},y_{2},\sigma_{2})\right|\leq L_{z,1}^{\star}\left|z_{1}-z_{2}\right|+L_{y,1}^{\star}\left|y_{1}-y_{2}\right|+L_{\sigma,1}^{\star}\left|\sigma_{1}-\sigma_{2}\right|,$
where the Lipschitz constants satisfy:
$L_{z,1}^{\star}$ $\leq\mu(Y_{LB})^{\gamma-2}(x^{\max}+Y_{UB})+M_{c^{\prime}},$
$L_{y,1}^{\star}$ $\leq 2\mu x^{\max}(Y_{LB})^{\gamma-2},$
$L_{\sigma,1}^{\star}$ $\leq 2\mu x^{\max}(Y_{LB})^{\gamma-1}.$
Proof of Lemma 18.
@@
Lipschitz continuity with respect to $z$.
@@
$\left|\frac{\partial H}{\partial z}(z,y,\sigma)\right|\leq\mu(Y_{LB})^{\gamma-2}(Y_{UB}+x^{\max})+M_{c^{\prime}}\eqqcolon L_{z,1}^{\star}.$
Lipschitz continuity with respect to $y$.
@@
Adding these,
$L_{y,1}^{\star}\leq 2\mu x^{\max}(Y_{LB})^{\gamma-2}.$
Lipschitz continuity with respect to $\sigma$.
@@
Summing,
$L_{\sigma,1}^{\star}\leq 2\mu x^{\max}(Y_{LB})^{\gamma-1}.$
Conclusion.
The function $\phi^{1}(z,y,\sigma)$ is Lipschitz continuous in each variable on its domain, and thus jointly Lipschitz. For any two points,
$\left|\phi^{1}(z_{1},y_{1},\sigma_{1})-\phi^{1}(z_{2},y_{2},\sigma_{2})\right|\leq L_{z,1}^{\star}\left|z_{1}-z_{2}\right|+L_{y,1}^{\star}\left|y_{1}-y_{2}\right|+L_{\sigma,1}^{\star}\left|\sigma_{1}-\sigma_{2}\right|,$
with the constants as given above. This completes the proof of Lemma 18
@@
Let $\phi_{k}(\mathbf{z},\sigma)$ be the maximum utility gain player $k$ can achieve by unilaterally deviating from the profile $(\mathbf{z},\mathbf{1})$ when the platform’s revenue sharing parameter is $\sigma$. For any profiles $\mathbf{x},\mathbf{x}^{\prime}$ such that $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$ and $Y_{LB}\leq\left\|\mathbf{x}_{-k}\right\|,\left\|\mathbf{x}^{\prime}_{-k}\right\|\leq Y_{UB}$, and any $\rho_{1},\rho_{2}$ such that $\left|\rho_{1}-\rho_{2}\right|\leq\delta$, we have:
$\left|\phi_{k}(\mathbf{x},\rho_{1})-\phi_{k}(\mathbf{x}^{\prime},\rho_{2})\right|\leq\delta\cdot\left(M_{c^{\prime}}+\mu\cdot B(Y_{LB},Y_{UB},n)\right),$
where $M_{c^{\prime}}=\max_{j}\sup_{x\in[0,x^{\max}]}\left|c_{j}^{\prime}(x)\right|$, $x^{\max}=\max_{j}x_{j}^{\max}$, and
$B(Y_{LB},Y_{UB},n)=(Y_{LB})^{\gamma-3}\Big{[}x^{\max}Y_{UB}(1+(4+\alpha)\sqrt{n}+2Y_{LB})+Y_{UB}^{2}+\sqrt{n}(1+\alpha)(x^{\max})^{2}\Big{]}.$
Proof of Lemma 19.
@@
Here, $x^{\max}=\max_{j}x_{j}^{\max}$ represents an upper bound on rational quality choices as established in Lemma 10.
To establish the Lipschitz continuity of $\phi_{k}$, we need to bound how much $\phi_{k}$ can change when its arguments vary within specified ranges. We’ll derive these bounds using the results from Lemmas 17 and 18, which provide upper bounds on the Lipschitz constants of the marginals of $\phi_{k}^{(s^{\prime}=0)}$ and $\phi_{k}^{(s^{\prime}=1)}$.
Step 1: Establishing Lipschitz constants for $\phi_{k}^{(s^{\prime}=1)}(z_{k},y_{k},\sigma)$.
Using Lemma 18, we obtain the following bounds on the Lipschitz constants:
$L_{z,1}$ $=\mu\cdot\left((Y_{LB})^{\gamma-2}(x^{\max}+Y_{UB})\right)+M_{c^{\prime}},$
$L_{y,1}$ $=2\mu x^{\max}(Y_{LB})^{\gamma-2},$
$L_{\sigma,1}$ $=2\mu x^{\max}(Y_{LB})^{\gamma-1}.$
Step 2: Establishing Lipschitz constants for $\phi_{k}^{(s^{\prime}=0)}(z_{k},y_{k},\sigma)$.
Similarly Lemma 17 ensures that:
$L_{z,0}$ $=\mu\cdot\left((Y_{LB})^{\gamma-2}(x^{\max}+Y_{UB})\right)+M_{c^{\prime}},$
$L_{y,0}$ $=\mu x^{\max}(Y_{LB})^{\gamma-2}+\frac{\mu\,x^{\max}(1+\alpha)}{(Y_{LB})^{3-\gamma}}\left[x^{\max}+Y_{UB}\right],$
$L_{\sigma,0}$ $=\mu x^{\max}(Y_{LB})^{\gamma-1}.$
Step 3: Lipschitz constants for $\phi_{k}(z_{k},y_{k},\sigma)$.
For $\phi_{k}(z_{k},y_{k},\sigma)=\max(\phi_{k}^{(s^{\prime}=0)},\phi_{k}^{(s^{\prime}=1)})$, the Lipschitz constants can be derived as follows:
If functions $f_{A}$ and $f_{B}$ are Lipschitz continuous with constants $L_{A}$ and $L_{B}$ respectively, then their maximum $\max(f_{A},f_{B})$ is Lipschitz continuous with constant $\max(L_{A},L_{B})$. Applying this principle to each of the three arguments of $\phi_{k}$, we have:
$L_{\phi,z_{k}}$ $=\max(L_{z,0},L_{z,1})=\mu\cdot\left((Y_{LB})^{\gamma-2}(x^{\max}+Y_{UB})\right)+M_{c^{\prime}},$
$L_{\phi,y_{k}}$ $=\max(L_{y,0},L_{y,1})$
$\leq\mu x^{\max}(Y_{LB})^{\gamma-2}+\frac{\mu x^{\max}(1+\alpha)}{(Y_{LB})^{3-\gamma}}\left[x^{\max}+Y_{UB}\right]+2\mu x^{\max}(Y_{LB})^{\gamma-2}$
$=3\mu x^{\max}(Y_{LB})^{\gamma-2}+\frac{\mu x^{\max}(1+\alpha)}{(Y_{LB})^{3-\gamma}}\left[x^{\max}+Y_{UB}\right],$
$L_{\phi,\sigma}$ $=\max(L_{\sigma,0},L_{\sigma,1})=\max\left(\mu x^{\max}(Y_{LB})^{\gamma-1},2\mu x^{\max}(Y_{LB})^{\gamma-1}\right)$
$=2\mu x^{\max}(Y_{LB})^{\gamma-1}.$
Step 4: Bounding the difference $\left|\phi_{k}(\mathbf{x},\rho_{1})-\phi_{k}(\mathbf{x}^{\prime},\rho_{2})\right|$.
@@
Using the Lipschitz constants derived above and the mean value theorem, we can bound the difference:
$\left|\phi_{k}(\mathbf{x},\rho_{1})-\phi_{k}(\mathbf{x}^{\prime},\rho_{2})\right|$ $\leq L_{\phi,z_{k}}\left|x_{k}-x^{\prime}_{k}\right|+L_{\phi,y_{k}}\left|y_{k}-y^{\prime}_{k}\right|+L_{\phi,\sigma}\left|\rho_{1}-\rho_{2}\right|$
$\leq L_{\phi,z_{k}}\delta+L_{\phi,y_{k}}\sqrt{n}\delta+L_{\phi,\sigma}\delta$
$=\delta\left(L_{\phi,z_{k}}+L_{\phi,y_{k}}\sqrt{n}+L_{\phi,\sigma}\right).$
Step 5: Combining the bounds.
Substituting our Lipschitz constants:
$L_{\phi,z_{k}}+L_{\phi,y_{k}}\sqrt{n}+L_{\phi,\sigma}$ $=\mu\cdot\left((Y_{LB})^{\gamma-2}(x^{\max}+Y_{UB})\right)+M_{c^{\prime}}$
@@
$+\sqrt{n}\left(3\mu x^{\max}(Y_{LB})^{\gamma-2}+\frac{\mu\,x^{\max}(1+\alpha)}{(Y_{LB})^{3-\gamma}}\left[x^{\max}+Y_{UB}\right]\right)$
$+2\mu x^{\max}(Y_{LB})^{\gamma-1}.$
Using $Y_{LB}\leq Y_{UB}$ and some manipulations, we can further bound the expression:
$L_{\phi,z_{k}}+L_{\phi,y_{k}}\sqrt{n}+L_{\phi,\sigma}$ $\leq\mu\cdot\left((Y_{LB})^{\gamma-3}Y_{UB}(x^{\max}+Y_{UB})\right)+M_{c^{\prime}}$
$+\sqrt{n}\Bigl{[}3\mu x^{\max}(Y_{LB})^{\gamma-3}Y_{UB}$
$+\mu\,x^{\max}(Y_{LB})^{\gamma-3}(1+\alpha)\left[x^{\max}+Y_{UB}\right]\Bigr{]}$
$+2\mu x^{\max}(Y_{LB})^{\gamma-3}Y_{LB}Y_{UB}$
$\leq M_{c^{\prime}}+\mu(Y_{LB})^{\gamma-3}\Big{[}x^{\max}Y_{UB}(1+(4+\alpha)\sqrt{n}+2Y_{LB})$
$+Y_{UB}^{2}+\sqrt{n}(1+\alpha)(x^{\max})^{2}\Big{]}.$
Rearranging:
$\leq M_{c^{\prime}}+\mu(Y_{LB})^{\gamma-3}\Big{[}x^{\max}Y_{UB}(1+(4+\alpha)\sqrt{n}+2Y_{LB})+Y_{UB}^{2}+\sqrt{n}(1+\alpha)(x^{\max})^{2}\Big{]}.$
This completes the proof of Lemma 19. ∎
Proposition 20.
Let $\rho_{1},\rho_{2}\in[0,1]$ such that $\left|\rho_{1}-\rho_{2}\right|\leq\delta$, and $\mathbf{x},\mathbf{x}^{\prime}$ such that $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$ and one of those profiles is an ESE in either $G(\rho_{1})$ or $G(\rho_{2})$. Let $L$ be the Lipschitz constant from Proposition 16, and define $Y_{LB}=\min_{k\in[n]}\sum_{i\neq k}x_{i}^{\min}$. Assuming $0.5Y_{LB}>\sqrt{n}\delta$, if $\mathbf{x}$ is an $\eta$-FSE in $G(\rho_{2})$, then $\mathbf{x}^{\prime}$ is a $(\eta+M_{1}\delta)$-FSE in $G(\rho_{1})$, where $M_{1}$ is a constant defined in the proof.
Proof of Proposition 20.
Without loss of generality, we assume $\mathbf{x}^{\prime}$ is the unique ESE in $G(\rho_{1})$, and denote it for the rest of the proof as $\mathcal{X}(\rho_{1})$.
Let $\phi_{k}(\mathbf{z},\sigma)$ denote the maximum utility gain player $k$ can achieve by unilaterally deviating from the profile $(\mathbf{z},\mathbf{1})$ when the platform’s revenue sharing parameter is $\sigma$. To apply Lemma 19, we need to establish bounds on other players’ contributions.
From Lemma 14, each player’s quality at ESE is bounded from below by their respective minimum quality $x_{i}^{\min}$. Thus, a lower bound on the total quality of other players for each player $k$ is given by $Y_{LB,k}=\sum_{i\neq k}x_{i}^{\min}$. We defined $Y_{LB}=\min_{k\in[n]}Y_{LB,k}$ in the proposition statement, which provides a valid lower bound on $\left\|\mathcal{X}(\rho_{1})_{-k}\right\|$ for any player $k$ and any $\rho\in[0,1]$. For the upper bound, we use the fact that each player’s quality is bounded by their respective $x_{i}^{\max}$ (from Lemma 10) to get $Y_{UB}=\sum_{j=1}^{n}x_{j}^{\max}-\min_{k}x_{k}^{\max}$. This gives us $Y_{LB}\leq\left\|\mathcal{X}(\rho_{1})_{-k}\right\|\leq Y_{UB}$ for any player $k$.
By assuming that $0.5Y_{LB}>\sqrt{n}\delta$, we ensure that the variations in equilibrium profiles do not violate the lower bounds needed for applying Lemma 19. Specifically, for any player $k$:
$\left|\left\|\mathcal{X}(\rho_{1})_{-k}\right\|-\left\|\mathbf{x}_{-k}\right\|\right|\leq\left\|\mathcal{X}(\rho_{1})_{-k}-\mathbf{x}_{-k}\right\|\leq\sqrt{n-1}\left\|\mathcal{X}(\rho_{1})_{-k}-\mathbf{x}_{-k}\right\|_{2}\leq\sqrt{n}\delta.$
Since $\left\|\mathcal{X}(\rho_{1})_{-k}\right\|\geq Y_{LB}$ and $0.5Y_{LB}\geq\sqrt{n}\delta$, we have $\left\|\mathbf{x}_{-k}\right\|\geq Y_{LB}-\sqrt{n}\delta>0.5Y_{LB}>0$, ensuring the bounds remain valid. Additionally, we have $\left\|\mathbf{x}_{-k}\right\|\leq Y_{UB}+\sqrt{n}\delta<1.5Y_{UB}$, which is also valid.
Now, we apply Lemma 19 with those bounds. The lemma states that for any profiles $\mathbf{z},\mathbf{z}^{\prime}$ with $\left\|\mathbf{z}-\mathbf{z}^{\prime}\right\|_{2}\leq\delta$ and revenue shares $\sigma_{1},\sigma_{2}$ with $\left|\sigma_{1}-\sigma_{2}\right|\leq\delta$:
@@
$\left|\phi_{k}(\mathbf{z},\sigma_{1})-\phi_{k}(\mathbf{z}^{\prime},\sigma_{2})\right|\leq\delta\cdot\left(M_{c^{\prime}}+\mu\cdot B(Y_{LB},Y_{UB},n)\right).$
where $M_{c^{\prime}}=\max_{j}\sup_{x\in[0,x_{j}^{\max}]}\left|c_{j}^{\prime}(x)\right|$ and
$B(Y_{LB},Y_{UB},n)=(Y_{LB})^{\gamma-3}\Big{[}x^{\max}Y_{UB}(1+(4+\alpha)\sqrt{n}+2Y_{LB})+Y_{UB}^{2}+\sqrt{n}(1+\alpha)(x^{\max})^{2}\Big{]}.$
By the assumption that $\mathbf{x}$ is an $\eta$-FSE in $G(\rho_{2})$, we know that $\phi_{k}(\mathbf{x},\rho_{2})\leq\eta$ for all players $k$.
Taking $\mathbf{z}=\mathcal{X}(\rho_{1}),\mathbf{z}^{\prime}=\mathbf{x},\sigma_{1}=\rho_{1},\sigma_{2}=\rho_{2}$, and applying the triangle inequality:
$\phi_{k}(\mathcal{X}(\rho_{1}),\rho_{1})\leq\left|\phi_{k}(\mathcal{X}(\rho_{1}),\rho_{1})-\phi_{k}(\mathbf{x},\rho_{2})\right|+\phi_{k}(\mathbf{x},\rho_{2}).$
Since $\left\|\mathbf{x}-\mathcal{X}(\rho_{1})\right\|_{2}\leq\delta$ and $\left|\rho_{1}-\rho_{2}\right|\leq\delta$, we have:
$\phi_{k}(\mathcal{X}(\rho_{1}),\rho_{1})\leq\delta\cdot\left(M_{c^{\prime}}+\mu\cdot B(0.5Y_{LB},1.5Y_{UB},n)\right)+\eta.$
Setting $M_{1}=M_{c^{\prime}}+\mu\cdot B(0.5Y_{LB},1.5Y_{UB},n)$, we get:
$\phi_{k}(\mathcal{X}(\rho_{1}),\rho_{1})\leq M_{1}\cdot\delta+\eta.$
This holds for all players $k$, which means that $\mathcal{X}(\rho_{1})$ is a $(\eta+M_{1}\delta)$-FSE in $G(\rho_{1})$, completing the proof of Proposition 20. ∎
@@
Fix some $\delta>0$. Suppose we have two quality profiles $\mathbf{x}$ and $\mathbf{x}^{\prime}$, and two revenue sharing parameters $\rho$ and $\rho^{\prime}$, such that the $\ell_{1}$-norms of the quality profiles are bounded: $\left\|\mathbf{x}\right\|_{1},\left\|\mathbf{x}^{\prime}\right\|_{1}\in[X_{LB},X_{UB}]$ for some fixed bounds $X_{LB}>0$ and $X_{UB}$, the $\ell_{2}$ distance between profiles is bounded: $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$, and the difference in revenue sharing parameters is $\left|\rho-\rho^{\prime}\right|\leq\delta$. Then, the absolute difference in the platform’s revenue is bounded by:
$\left|U_{P}(\mathbf{x},\mathbf{1};f_{\rho})-U_{P}(\mathbf{x}^{\prime},\mathbf{1};f_{\rho^{\prime}})\right|\leq D(X_{LB},X_{UB},n,\gamma,\mu)\cdot\delta,$
where $D(X_{LB},X_{UB},n,\gamma,\mu)=\mu\left(\gamma(X_{LB})^{\gamma-1}\sqrt{n}+(X_{UB})^{\gamma}\right)$.
Proof of Lemma 21.
Under full sharing ($\mathbf{s}=\mathbf{1}$), the platform’s revenue function is given by $U_{P}(\mathbf{x},\mathbf{1};f_{\rho})=\frac{\mu\alpha}{1+\alpha}(1-\rho)\left\|\mathbf{x}\right\|_{1}^{\gamma}$. This follows from the definition of the platform’s revenue and the proportional allocation rule: The GenAI quality is $\alpha\left\|\mathbf{x}\right\|$, the platform’s traffic is $\mu\left\|\mathbf{x}\right\|^{\gamma}$, and the platform retains a fraction $(1-\rho)$ of the AI-driven revenue.
To bound the difference in utilities, we apply the triangle inequality to decompose it into two components:
@@
$=C\left|1-\rho\right|\cdot\left|\left\|\mathbf{x}\right\|^{\gamma}-\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\right|+C\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\cdot\left|\rho^{\prime}-\rho\right|.$
For the first term, since $\rho\in[0,1]$, we have $\left|1-\rho\right|\leq 1$. To bound the difference of powers $\left|\left\|\mathbf{x}\right\|^{\gamma}-\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\right|$, we use the Mean Value Theorem. Let $g(v)=v^{\gamma}$ for $v>0$, where $\gamma\in[0,1]$. By the Mean Value Theorem, there exists a value $\xi$ between $\left\|\mathbf{x}\right\|$ and $\left\|\mathbf{x}^{\prime}\right\|$ such that:
$\left|\left\|\mathbf{x}\right\|^{\gamma}-\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\right|=\left|g(\left\|\mathbf{x}\right\|)-g(\left\|\mathbf{x}^{\prime}\right\|)\right|=\left|g^{\prime}(\xi)\right|\cdot\left|\left\|\mathbf{x}\right\|-\left\|\mathbf{x}^{\prime}\right\|\right|.$
@@
$\left|\left\|\mathbf{x}\right\|^{\gamma}-\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\right|\leq\gamma(X_{LB})^{\gamma-1}\cdot\left|\left\|\mathbf{x}\right\|-\left\|\mathbf{x}^{\prime}\right\|\right|.$
@@
$=D(X_{LB},X_{UB},n,\gamma,\mu)\cdot\delta,$
where $D(X_{LB},X_{UB},n,\gamma,\mu)=\mu\left(\gamma(X_{LB})^{\gamma-1}\sqrt{n}+(X_{UB})^{\gamma}\right)$ as defined in the lemma statement. This completes the proof of Lemma 21. ∎
Proposition 22.
Let $\rho_{1},\rho_{2}\in[0,1]$ such that $\left|\rho_{1}-\rho_{2}\right|\leq\delta$, and $\mathbf{x},\mathbf{x}^{\prime}$ such that $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$ and one of those profiles is an ESE in either $G(\rho_{1})$ or $G(\rho_{2})$. Let $X_{LB}=\sum_{i=1}^{n}x_{i}^{\min}$ and $X_{UB}=\sum_{j=1}^{n}x_{j}^{\max}$. Assuming $0.5X_{LB}\geq\sqrt{n}\delta$, it holds that
$\left|U_{P}(\mathcal{X}(\rho_{1}),\mathbf{1};f_{\rho_{1}})-U_{P}(\mathbf{x},\mathbf{1};f_{\rho_{2}})\right|\leq M_{2}\delta,$
where $M_{2}$ is a constant defined in the proof.
Proof of Proposition 22.
Without loss of generality, we assume $\mathbf{x}^{\prime}$ is the unique ESE in $G(\rho_{1})$, and denote it for the rest of the proof as $\mathcal{X}(\rho_{1})$. We need to show that the difference in platform revenues is bounded by a constant times $\delta$.
First, we verify that the quality profiles $\mathcal{X}(\rho_{1})$ and $\mathbf{x}$ satisfy the conditions required by Lemma 21. Specifically, we need to establish appropriate bounds on their $\ell_{1}$-norms.
From Lemmas 14 and 10, we know that the following bounds hold for $\left\|\mathcal{X}(\rho_{1})\right\|$:
$X_{LB}\leq\left\|\mathcal{X}(\rho_{1})\right\|\leq X_{UB}.$
For the profile $\mathbf{x}$, we can derive bounds using the proximity to $\mathcal{X}(\rho_{1})$. By the reverse triangle inequality and the relationship between the norms, we have:
$\left|\left\|\mathcal{X}(\rho_{1})\right\|-\left\|\mathbf{x}\right\|\right|\leq\left\|\mathcal{X}(\rho_{1})-\mathbf{x}\right\|\leq\sqrt{n}\left\|\mathcal{X}(\rho_{1})-\mathbf{x}\right\|_{2}\leq\sqrt{n}\delta.$
This implies:
$\left\|\mathcal{X}(\rho_{1})\right\|-\sqrt{n}\delta\leq\left\|\mathbf{x}\right\|\leq\left\|\mathcal{X}(\rho_{1})\right\|+\sqrt{n}\delta.$
Combining this with our bounds on $\left\|\mathcal{X}(\rho_{1})\right\|$, we obtain:
$X_{LB}-\sqrt{n}\delta\leq\left\|\mathbf{x}\right\|\leq X_{UB}+\sqrt{n}\delta.$
Given our assumption that $0.5X_{LB}\geq\sqrt{n}\delta$, we can ensure that $X_{LB}-\sqrt{n}\delta\geq 0.5X_{LB}>0$, which guarantees that the lower bound remains positive - a necessary condition for applying the lemma.
Thus, we can define:
$X_{LB}=X_{LB}-\sqrt{n}\delta\geq 0.5X_{LB}>0,$
$X_{UB}=X_{UB}+\sqrt{n}\delta\leq 1.5X_{UB}.$
Now, we can apply Lemma 21 with these bounds. The lemma states:
$\left|U_{P}(\mathcal{X}(\rho_{1}),\mathbf{1};f_{\rho_{1}})-U_{P}(\mathbf{x},\mathbf{1};f_{\rho_{2}})\right|\leq D(X_{LB},X_{UB},n,\gamma,\mu)\cdot\delta,$
where
$D(X_{LB},X_{UB},n,\gamma,\mu)=\mu\left(\gamma(X_{LB})^{\gamma-1}\sqrt{n}+(X_{UB})^{\gamma}\right).$
Substituting our bounds $Y_{LB}\geq 0.5X_{LB}$ and $X_{UB}\leq 1.5X_{UB}$, we have:
$D(Y_{LB},Y_{UB},n,\gamma,\mu)\leq\mu\left(\gamma(0.5X_{LB})^{\gamma-1}\sqrt{n}+(1.5X_{UB})^{\gamma}\right).$
Therefore:
$\left|U_{P}(\mathcal{X}(\rho_{1}),\mathbf{1};f_{\rho_{1}})-U_{P}(\mathbf{x},\mathbf{1};f_{\rho_{2}})\right|\leq\delta\cdot\mu\left(\gamma(0.5X_{LB})^{\gamma-1}\sqrt{n}+(1.5X_{UB})^{\gamma}\right).$
Defining $M_{2}=\mu\left(\gamma(0.5X_{LB})^{\gamma-1}\sqrt{n}+(1.5X_{UB})^{\gamma}\right)$ completes the proof of Proposition 22. ∎

### v2

@@ Conclusion
Define $K(\sigma)\coloneqq\frac{\mu(1+\alpha\sigma)}{1+\alpha}$, and define for any $(z,y,\sigma)\in[0,x^{\max}]\times[Y_{LB},Y_{UB}]\times[0,1]$:
@@
$\left|\phi^{1}(z_{1},y_{1},\sigma_{1})-\phi^{1}(z_{2},y_{2},\sigma_{2})\right|$
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2,0,1,2-\gamma)\left|y_{1}-y_{2}\right|+B(0,2,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|$
Proof of Lemma˜21.
@@
Lipschitz continuity with respect to $z$
@@
$\left|\frac{\partial H}{\partial z}(z,y,\sigma)\right|\leq\mu(Y_{LB})^{\gamma-2}(Y_{UB}+x^{\max})+M_{c^{\prime}}\leq B(1,1,0,1,2-\gamma).$
Lipschitz continuity with respect to $y$
@@
Adding these, we get:
$2\mu x^{\max}(Y_{LB})^{\gamma-2}\leq B(0,2,0,1,2-\gamma).$
Lipschitz continuity with respect to $\sigma$
@@
Summing, we get:
$2\mu x^{\max}(Y_{LB})^{\gamma-1}\leq B(0,2,0,1,1-\gamma).$
Conclusion
Combining the three Lipschitz constants, we have:
$\left|\phi^{1}(z_{1},y_{1},\sigma_{1})-\phi^{1}(z_{2},y_{2},\sigma_{2})\right|$
$\leq B(1,1,0,1,2-\gamma)\left|z_{1}-z_{2}\right|+B(0,2,0,1,2-\gamma)\left|y_{1}-y_{2}\right|+B(0,2,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|$
This completes the proof of Lemma˜21
@@
Let $\phi_{k}(\mathbf{z},\sigma)$ be the maximum utility gain player $k$ can achieve by unilaterally deviating from the profile $(\mathbf{z},\mathbf{1})$ when the platform’s revenue sharing parameter is $\sigma$. For any profiles $\mathbf{x},\mathbf{x}^{\prime}$ such that $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$ and $Y_{LB}\leq\left\|\mathbf{x}_{-k}\right\|,\left\|\mathbf{x}^{\prime}_{-k}\right\|\leq Y_{UB}$, and any $\rho,\rho^{\prime}$ such that $\left|\rho-\rho^{\prime}\right|\leq\delta$, we have:
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq\delta\left(B(1,1,0,1,2-\gamma)+B(0,(2+\alpha),0.5,2,3-\gamma)+B(0,2,0,1,1-\gamma)\right)$
Proof of Lemma˜22.
@@
Step 1: Lipschitz continuity of deviation gain function $\phi_{k}$
@@
Recall that the Lipschitz constant of a maximum of two functions is at most the maximum of the Lipschitz constants of the two functions. Therefore, we can bound the Lipschitz constant of $\phi_{k}$ by taking the maximum of the Lipschitz constants of $\phi_{k}^{(s^{\prime}=0)}$ and $\phi_{k}^{(s^{\prime}=1)}$:
  • For the $z$ coordinate:
    $\max\{B(1,1,0,1,2-\gamma),B(1,1,0,1,2-\gamma)\}=B(1,1,0,1,2-\gamma).$
  • For the $y$ coordinate:
    $\max\{B(0,2+\alpha,0,2,3-\gamma),B(0,2,0,1,2-\gamma)\}=B(0,2+\alpha,0,2,3-\gamma).$
  • For the $\sigma$ coordinate:
    $\max\{B(0,1,0,1,1-\gamma),B(0,2,0,1,1-\gamma)\}=B(0,2,0,1,1-\gamma).$
Writing it explicitly, we have:
$\left|\phi_{k}(z_{k,1},y_{k,1},\sigma_{1})-\phi_{k}(z_{k,2},y_{k,2},\sigma_{2})\right|$
$\leq B(1,1,0,1,2-\gamma)\left|z_{k,1}-z_{k,2}\right|+B(0,2+\alpha,0,2,3-\gamma)\left|y_{k,1}-y_{k,2}\right|+B(0,2,0,1,1-\gamma)\left|\sigma_{1}-\sigma_{2}\right|.$
Step 2: Bounding the difference $\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|$
@@
Using the Lipschitz constant derived above and the mean value theorem, we can bound the difference:
@@
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq B(1,1,0,1,2-\gamma)\delta+B(0,2+\alpha,0,2,3-\gamma)\sqrt{n}\delta+B(0,2,0,1,1-\gamma)\delta\\$
Notice we can write the expression $B(0,2+\alpha,0,2,3-\gamma)\sqrt{n}$ as $B(0,(2+\alpha),0.5,2,3-\gamma)$, as the third parameter affects the exponent of $n$ in the definition of $B$. Thus, we can combine the terms to get:
@@
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq\delta\cdot\left(B(1,1,0,1,2-\gamma)+B(0,(2+\alpha),0.5,2,3-\gamma)+B(0,2,0,1,1-\gamma)\right)$
@@
Fix some $\delta>0$. Let two quality profiles $\mathbf{x}$ and $\mathbf{x}^{\prime}$ such that $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$, and two revenue sharing parameters $\rho,\rho^{\prime}$ such that $\left|\rho-\rho^{\prime}\right|\leq\delta$. Then, the absolute difference in the platform’s revenue is bounded by:
$\left|U_{P}(\mathbf{x},\mathbf{1};f_{\rho})-U_{P}(\mathbf{x}^{\prime},\mathbf{1};f_{\rho^{\prime}})\right|\leq B(0,1+\gamma,0.5,1,1-\gamma)\cdot\delta.$
Proof of Lemma˜23.
We apply the triangle inequality to decompose the difference in utilities into two components:
@@
$=C(1-\rho)\cdot\left|\left\|\mathbf{x}\right\|^{\gamma}-\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\right|+C\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\cdot\left|\rho^{\prime}-\rho\right|.$
To bound the difference of powers $\left|\left\|\mathbf{x}\right\|^{\gamma}-\left\|\mathbf{x}^{\prime}\right\|^{\gamma}\right|$, we use the Mean Value Theorem. Let $g(v)=v^{\gamma}$ for $v>0$, where $\gamma\in[0,1]$. By the Mean Value Theorem, there exists a value $\xi$ between $\left\|\mathbf{x}\right\|$ and $\left\|\mathbf{x}^{\prime}\right\|$ such that:
[…]
@@
[…]
@@
$\leq\left(B(0,\gamma,0.5,0,1-\gamma)+B(0,1,0,\gamma,0)\right)\cdot\delta$
$\leq B(0,1+\gamma,0.5,1,1-\gamma)\cdot\delta.$
This completes the proof of Lemma˜23. ∎
C.3 Lipschitz Continuity Results Used in the Proof of Algorithm˜1